40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 764     Entry time: Wed Jul 30 12:03:44 2008
Author: Masha 
Type: Summary 
Category: Auxiliary locking 
Subject: weekly summary 
I've been learning about mode matching/beam propagation, so I can work on getting more
light into the fiber and increase the phase noise signal. I am also looking into phase
lock loops and noise in the fiber stabilization system to understand the noise sources
and figure out what our goals are in fiber stabilization.

In the lab, I've reproduced the Mach Zehnder interferometer that I had at the 40m, now
with a 50m fiber in one arm. I have done some preliminary fiber noise measurements
and revised estimations of noise sources (see attached plots). Once the digital
acquisition system is back up, I will be able to better manipulate the signals to cancel
laser amplitude noise and amplitude noise from variation in the amount of coupling into
the fiber. Some improvements in progress are more stable mounts for the fiber couplings,
faraday isolator, and better mode matching with the fiber.

Also working on my progress report.
Attachment 1: mz_fiber_noise0730.png  99 kB  | Show | Hide all | Show all
Attachment 2: setup0730.JPG  93 kB  | Hide | Hide all | Show all
setup0730.JPG
Attachment 3: MZnoise_sources0727.png  30 kB  Uploaded Wed Jul 30 13:07:33 2008  | Hide | Hide all | Show all
MZnoise_sources0727.png
ELOG V3.1.3-